OpenStudy (anonymous):

Find the x-, y, and z-intercepts for the equation. 2x - 6y - 3z = 12

OpenStudy (anonymous):

for x intercept, y and z are 0

OpenStudy (anonymous):

2x - 6y - 3z = 12 To find the x-intercept, take y and z as 0. 2x - 0 - 0 = 12 2x = 12 x = 12/2 = 6 To find the y-intercept, take x and z as 0. 0 - 6y - 0 = 12 -6y = 12 -12 = 6y -12/6 = -2 = y To find the z-intercept, take y and x as 0. 0 + 0 - 3z = 12 -12 = 3z -12/3 = - 4 = z

OpenStudy (anonymous):

Thanks Guys :)